OpenStudy (anonymous):

convert to polar form x=2

OpenStudy (turingtest):

same formula as I showed you in the last question did you not understand it?

OpenStudy (anonymous):

not really, I was sick when my class went over this so I have no idea what I'm doing

OpenStudy (turingtest):

look at your other post I just responded to

OpenStudy (anonymous):

So for polar form, we will give the position as (r, theta). Since x=2 is a vertical line, we get the radius varying with theta.

OpenStudy (anonymous):

It looks to me that r=2sec(theta)

OpenStudy (anonymous):

So your answer in polar form is \[(2\sec \theta,\theta)~~-\frac{\pi}{2}<\theta<\frac{\pi}{2}\]